OpenStudy (anonymous):

write the standard form of the equation of the line that passes through B(-10,-5) and is perpendicular to the graph of 6x-5y=24

OpenStudy (anonymous):

line perpendicular to 6x-5y=24 is of the form 5x+6y=k since it passes through (-10,-5) hence 5(-10) +6(-5)=k or k=-80 hence the reqd equation is 5x+6y=-80 or 5x+6y+80=0